OpenStudy (anonymous):

Find the domain of the function g(x)=6/2-7x

OpenStudy (anonymous):

All real numbers except x=2/7

hero (hero):

\[x \neq \frac {2} {7}\]

OpenStudy (anonymous):

{x|x is a real number and x > = 2/7}?

OpenStudy (anonymous):

why does x need to be larger then 2/7? the function can be negative valued.

OpenStudy (anonymous):

or equal to

OpenStudy (anonymous):

there are four choices to pick

OpenStudy (anonymous):

that was the first

OpenStudy (anonymous):

i am not familiar with the equal sign with a slash through it

hero (hero):

\[(-\infty, \frac {2}{7}), (\frac {2}{7}, \infty)\]

hero (hero):

the slash means not equal

hero (hero):

What are the other choices?

OpenStudy (anonymous):

{x|x is a real number and x is not equal to 2/7}?

hero (hero):

Yes, that's it